# Combinatorial interpretation of a truncated identity of Gauss

### v1

@@
Abstract. In 2012, Andrews and Merca proved a truncated theorem on Euler’s pentagonal number theorem, which opened up a new study on truncated theta series. In particular, some results on a truncated identity of Guass have been proved. In this article, we provide a new combinatorial interpretation of the truncated identity of Gauss in terms of the minimal excludant non-overlined part of an overpartition.
Keywords: truncated identity of Gauss, the minimal excludant integer, overpartitions, non-overlined parts
@@ 1 Introduction
Apart from Euler’s pentagonal number theorem, there is one other classical theta identity due to Guass [1, p. 23]:
@@
An overpartition, introduced by Corteel and Lovejoy [7], is a partition such that the last occurrence of a number may be overlined. For example, there are fourteen overpartitions of 44.
@@
A combinatorial proof of (1.6) was given by Guo and Zeng [9]. One of the main objectives of this article is to present a new combinatorial proof of (1.6).
In 2018, Andrews and Merca [6] provided the following version of (1.3). For $k\geq 1$,
$\frac{(-q;q)_{\infty}}{(q;q)_{\infty}}\left(1+2\sum_{j=1}^{k}(-1)^{j}q^{j^{2}}\right)=1+2(-1)^{k}\frac{(-q;q)_{k}}{(q;q)_{k}}\sum_{j=0}^{\infty}\frac{q^{(k+1)(k+j+1)}(-q^{k+j+2};q)_{\infty}}{(1-q^{k+j+1})(q^{k+j+2};q)_{\infty}},$
@@
In this article, we will present a new combinatorial interpretation of the sum in (1.5) by using the minimal excludant non-overlined part of an overpartition. The minimal excludant of a partition was introduced by Grabner and Knopfmacher [8] under the name “smallest gap”. Recently, Andrews and Newman [3] undertook a combinatorial study of the minimal excludant of a partition. The minimal excludant of a partition $\pi$ is the smallest positive integer that is not a part of $\pi$, denoted $mex(\pi)$. There is a generalization of the minimal excludant of a partition. In [4], Andrews and Newman defined $mex_{A,a}(\pi)$ to be the smallest positive integer congruent to $a$ modulo AA that is not a part of $\pi$. In [10], Yang and Zhou extended the definition of $mex_{A,a}(\pi)$ to overpartitions.
@@
[10, Definition 4.1] Let $\overline{mes}_{A,a}(\pi)$ be the smallest integer congruent to $a$ modulo AA that is not a non-overlined part in the overpartition $\pi$. For $n\geq 1$, define $op_{A,a}(n)$ (resp. $\overline{op}_{A,a}(n)$) to be the number of overpartitions $\pi$ of $n$ such that $\overline{mes}_{A,a}(\pi)$ is congruent to $a$ (resp. $a+A$) modulo $2A$.
@@
For $n\geq 1$ and $k\geq 0$, define $op_{2,1}(n,k)$ to be the number of overpartitions $\pi$ of $n$ such that $\overline{mes}_{2,1}(\pi)\geq 2k+1$ and $\overline{mes}_{2,1}(\pi)\equiv 2k+1\pmod{4}$. By definition, we have
@@
We give the following combinatorial interpretation of the sum in (1.5).
@@
$(-1)^{k}\left(\overline{p}(n)+2\sum_{j=1}^{k}(-1)^{j}\overline{p}(n-j^{2})\right)=2op_{2,1}(n,k+1).$
The following truncated version of (1.2) immediately follows from Theorem 1.5.
@@
This article is organized as follows. In Section 2, we give a combinatorial proof of (1.6). Then, we show Theorem 1.5 in Section 3.
@@
For $n\geq 1$, let $\mathcal{A}(n)$ be the set of overpartitions of $n$ such that the smallest part is non-overlined. For an overpartition $\pi$ in $\mathcal{A}(n)$, if we change the smallest part to an overlined part, then we get an overpartition of $n$ such that the smallest part is overlined, and vice versa. This implies that the number of overpartitions in $\mathcal{A}(n)$ is $\overline{p}(n)/2$.
For $n\geq 1$, let $\mathcal{B}(n-1)$ be the set of overpartitions $\lambda$ of n1n-1 such that if $\overline{1}$ appears in $\lambda$ then the smallest part greater than 11 in $\lambda$ is greater than or equal to two plus the number of parts 11 in $\lambda$.
For $n\geq 1$, we set $\mathcal{C}(n-1)$ be the set of overpartitions of n1n-1 not in $\mathcal{B}(n-1)$, that is, $\mathcal{C}(n-1)$ is the set of overpartitions $\lambda$ of n1n-1 such that $\overline{1}$ appears in $\lambda$ and the smallest part greater than 11 in $\lambda$ is less than two plus the number of parts 11 in $\lambda$.
@@
In either case, we get an overpartition $\lambda$ in $\mathcal{B}(n)$. Obviously, the process above is reversible. Now, we have built a bijection between $\mathcal{A}(n)$ and $\mathcal{B}(n-1)$ for $n\geq 1$.
@@
Proof. For an overpartition $\pi$ of $n-j^{2}$, if we add the parts $1,3,\ldots,2j-1$ to $\pi$, then we get an overpartition $\lambda$ of $n$ with $\overline{mes}_{2,1}(\lambda)\geq 2j+1$, and vice versa. It implies that $\overline{p}(n-j^{2})$ equals the number of overpartitions $\lambda$ of $n$ with $\overline{mes}_{2,1}(\lambda)\geq 2j+1$.
On the other hand, it follows from definition that $op_{2,1}(n,j)+op_{2,1}(n,j+1)$ is also the number of overpartitions $\lambda$ of $n$ with $\overline{mes}_{2,1}(\lambda)\geq 2j+1$. This completes the proof.
@@
Proof of Theorem 1.5. By Lemmas 3.1 and 3.3, we have
@@
$=2op_{2,1}(n,k+1),$
and thus the proof is complete.

### v2

# Combinatorial interpretations of truncated versions of a identity of Gauss

@@
Abstract. In 2012, Andrews and Merca proved a truncated theorem on Euler’s pentagonal number theorem, which opened up a new study on truncated theta series. In particular, some truncated versions of a identity of Gauss have been proved. In this article, we provide new combinatorial interpretations of the truncated versions of the identity of Gauss in terms of the minimal excludant non-overlined part of an overpartition.
Keywords: truncated versions, a identity of Gauss, the minimal excludant integer, overpartitions, non-overlined parts
@@ 1 Introduction
Apart from Euler’s pentagonal number theorem, there is one other classical theta identity due to Gauss [1, p. 23]:
@@
An overpartition, introduced by Corteel and Lovejoy [8], is a partition such that the last occurrence of a number may be overlined. For example, there are fourteen overpartitions of $4$.
@@
A combinatorial proof of (1.6) was given by Guo and Zeng [11]. In 2018, Andrews and Merca [3] provided the following truncated version of (1.2). For $k\geq 1$,
$\frac{(-q;q)_{\infty}}{(q;q)_{\infty}}\left(1+2\sum_{j=1}^{k}(-1)^{j}q^{j^{2}}\right)=1+2(-1)^{k}\frac{(-q;q)_{k}}{(q;q)_{k}}\sum_{j=0}^{\infty}\frac{q^{(k+1)(k+j+1)}(-q^{k+j+2};q)_{\infty}}{(q^{k+j+1};q)_{\infty}},$
@@
The minimal excludant of a partition was introduced by Grabner and Knopfmacher [10] under the name “smallest gap”. Recently, Andrews and Newman [4] undertook a combinatorial study of the minimal excludant of a partition. The minimal excludant of a partition $\pi$ is the smallest positive integer that is not a part of $\pi$, denoted $mex(\pi)$. In [6], Ballantine and Merca proved some inequalities involving the minimal excludant of a partition. They also posed the following conjecture.
Conjecture 1.4.
[6, Conjecture 1] For $n,k\geq 1$,
$\sum_{j=-\infty}^{\infty}(-1)^{j}\overline{M}_{k}(n-j(3j-1))\geq 0,$
with strict inequality if $n\geq(k+1)^{2}$.
Conjecture 1.4 was settled by Yao in [19]. Furthermore, Yao showed that for $\ell\geq 1$,
$\sum_{n=0}^{\infty}\sum_{j=-\infty}^{\infty}(-1)^{j}\overline{M}_{k}(n-\ell j(3j-1)/2)q^{n}=2\frac{(q^{\ell};q^{\ell})_{\infty}}{(q;q)_{\infty}}\sum_{j=0}^{\infty}\frac{q^{(k+2j+1)^{2}}(1-q^{2k+4j+3})}{(q;q^{2})_{\infty}}.$ (1.7)
In [11], Guo and Zeng conjectured a stronger inequality than (1.5).
Conjecture 1.5.
[11, (6.4)] For $n,k\geq 1$,
$(-1)^{k-1}\left(\overline{p}(n)+2\sum_{j=1}^{k}(-1)^{j}\overline{p}(n-j^{2})\right)+\overline{p}(n-k^{2})\geq 0,$ (1.8)
with strict inequality if $n\geq k^{2}$.
Conjecture 1.5 was proved independently by Mao [13] and Yee [20] and then reconfirmed later by Wang and Yee [14]. In [17], Xia, Yee and Zhao proved the following inequality which implies (1.8).
$(-1)^{k-1}\left(\overline{p}(n)+2\sum_{j=1}^{k}(-1)^{j}\overline{p}(n-j^{2})\right)+\overline{p}(n-k(k+1))\geq 0.$
In [12], Li obtained the following truncated version of (1.2). For $k\geq 1$,
$\frac{(-q;q)_{\infty}}{(q;q)_{\infty}}\left(1+\sum_{j=-k}^{k-1}(-1)^{j}q^{j^{2}}\right)=1+(-1)^{k-1}\frac{(-q;q)_{k}}{(q;q)_{k-1}}\sum_{j=0}^{\infty}\frac{q^{k(k+j)}(-q^{k+j+1};q)_{\infty}}{(q^{k+j+1};q)_{\infty}}.$ (1.9)
Li [12] and Wang and Xiao [15, 16] presented the combinatorial interpretations of the sum in (1.8).
2 Main results
In this section, we will list the main results in article. One of the main objectives of this article is to present a new combinatorial proof of (1.6).
There is a generalization of the minimal excludant of a partition. In [5], Andrews and Newman defined $mex_{A,a}(\pi)$ to be the smallest positive integer congruent to $a$ modulo $A$ that is not a part of $\pi$. In [18], Yang and Zhou extended the definition of $mex_{A,a}(\pi)$ to overpartitions.
@@ 2 Main results
[18, Definition 4.1] Let $\overline{mex}_{A,a}(\pi)$ be the smallest integer congruent to $a$ modulo $A$ that is not a non-overlined part in the overpartition $\pi$. For $n\geq 1$, define $op_{A,a}(n)$ (resp. $\overline{op}_{A,a}(n)$) to be the number of overpartitions $\pi$ of $n$ such that $\overline{mex}_{A,a}(\pi)$ is congruent to $a$ (resp. $a+A$) modulo $2A$.
@@
For $n\geq 1$ and $k\geq 0$, define $op_{2,1}(n,k)$ to be the number of overpartitions $\pi$ of $n$ such that $\overline{mex}_{2,1}(\pi)\geq 2k+1$ and $\overline{mex}_{2,1}(\pi)\equiv 2k+1\pmod{4}$. By definition, we have
@@
In this article, we will present new combinatorial interpretations of the sum in (1.5) and (1.8) by using the minimal excludant non-overlined part of an overpartition.
@@
$(-1)^{k}\left(\overline{p}(n)+2\sum_{j=1}^{k}(-1)^{j}\overline{p}(n-j^{2})\right)=2op_{2,1}(n,k+1),$ (2.3)
and
$(-1)^{k-1}\left(\overline{p}(n)+2\sum_{j=1}^{k}(-1)^{j}\overline{p}(n-j^{2})\right)+\overline{p}(n-k^{2})=op_{2,1}(n,k)-op_{2,1}(n,k+1).$
For $k\geq 1$, the generating function of $op_{2,1}(n,k+1)$ is
$\begin{split}\sum_{n=1}^{\infty}op_{2,1}(n,k+1)&=(-q;q)_{\infty}\sum_{j=0}^{\infty}\frac{q^{1+3+\cdots+(2k+4j+1)}(1-q^{2k+4j+3})}{(q;q)_{\infty}}\\ &=\frac{(-q;q)_{\infty}}{(q;q)_{\infty}}\sum_{j=0}^{\infty}q^{(k+2j+1)^{2}}(1-q^{2k+4j+3}).\end{split}$ (2.4)
Then, the following truncated version of (1.2) immediately follows from (2.3) and (2.4).
Corollary 2.3.
For $k\geq 1$,
@@
Combining Theorems 1.3 and 2.2, we can get the following corollary.
Corollary 2.4.
For $n,k\geq 1$,
$op_{2,1}(n,k+1)=\frac{\overline{M}_{k}(n)}{2},$ (2.5)
and
$(-1)^{k-1}\left(\overline{p}(n)+2\sum_{j=1}^{k}(-1)^{j}\overline{p}(n-j^{2})\right)+\overline{p}(n-k^{2})=\frac{\overline{M}_{k-1}(n)}{2}-\frac{\overline{M}_{k}(n)}{2}.$ (2.6)
Then, we can get (1.7) by using (1.1) with $q\rightarrow q^{\ell}$, (2.4), (2.5), and Euler’s partition identity [9] (see also [1, (1.2.5)]):
$\frac{1}{(q;q^{2})_{\infty}}=(-q;q)_{\infty}.$
Appealing to (1.9) and (2.6), we can get following corollary.
Corollary 2.5.
For $k\geq 1$,
$\sum_{n=1}^{\infty}\left(\overline{M}_{k-1}(n)-\overline{M}_{k}(n)\right)q^{n}=2\frac{(-q;q)_{k}}{(q;q)_{k-1}}\sum_{j=0}^{\infty}\frac{q^{k(k+j)}(-q^{k+j+1};q)_{\infty}}{(q^{k+j+1};q)_{\infty}}.$
In this article, we will present an analytic proof of Corollary 2.5.
This article is organized as follows. We first give a combinatorial proof of (1.6) in Section 3. Then, we show Theorem 2.2 in Section 4. Finally, we give an analytic proof of Corollary 2.5 in Section 5.
@@
For $n\geq 1$, let $\mathcal{A}(n)$ be the set of overpartitions of $n$ such that the smallest part is non-overlined. For example, we have
$\mathcal{A}(4)=\{(4),(3,1),(\overline{3},1),(2,2),(2,1,1),(\overline{2},1,1),(1,1,1,1)\}.$
For an overpartition $\pi$ in $\mathcal{A}(n)$, if we change the smallest part to an overlined part, then we get an overpartition of $n$ such that the smallest part is overlined, and vice versa. This implies that the number of overpartitions in $\mathcal{A}(n)$ is $\overline{p}(n)/2$.
For $n\geq 1$, let $\mathcal{B}(n-1)$ be the set of overpartitions $\lambda$ of n1n-1 such that if $\overline{1}$ appears in $\lambda$ then the smallest part greater than $1$ in $\lambda$ is greater than or equal to two plus the number of parts $1$ in $\lambda$. For example, we have
$\mathcal{B}(3)=\{(3),(\overline{3}),(2,1),(\overline{2},1),(2,\overline{1}),(1,1,1),(1,1,\overline{1})\}.$
For $n\geq 1$, we set $\mathcal{C}(n-1)$ be the set of overpartitions of n1n-1 not in $\mathcal{B}(n-1)$, that is, $\mathcal{C}(n-1)$ is the set of overpartitions $\lambda$ of n1n-1 such that $\overline{1}$ appears in $\lambda$ and the smallest part greater than $1$ in $\lambda$ is less than two plus the number of parts $1$ in $\lambda$. For example, we have
$\mathcal{C}(3)=\{(\overline{2},\overline{1})\}.$
@@
In either case, we get an overpartition $\lambda$ in $\mathcal{B}(n-1)$. Obviously, the process above is reversible. Now, we have built a bijection between $\mathcal{A}(n)$ and $\mathcal{B}(n-1)$ for $n\geq 1$.
@@
Proof. For an overpartition $\pi$ of $n-j^{2}$, if we add the parts $1,3,\ldots,2j-1$ into $\pi$, then we get an overpartition $\lambda$ of $n$ with $\overline{mex}_{2,1}(\lambda)\geq 2j+1$, and vice versa. It implies that $\overline{p}(n-j^{2})$ equals the number of overpartitions $\lambda$ of $n$ with $\overline{mex}_{2,1}(\lambda)\geq 2j+1$.
On the other hand, it follows from definition that $op_{2,1}(n,j)+op_{2,1}(n,j+1)$ is also the number of overpartitions $\lambda$ of $n$ with $\overline{mex}_{2,1}(\lambda)\geq 2j+1$. This completes the proof.
@@
Proof of Theorem 2.2. It follows from Lemma 4.1 that
$\overline{p}(n-k^{2})=op_{2,1}(n,k)+op_{2,1}(n,k+1).$
So, we just need to show (2.3). By Lemmas 4.1 and 4.3, we have
@@
$=2op_{2,1}(n,k+1).$
We arrive at (2.3), and thus the proof is complete.
5 Analytic proof of Corollary 2.5
In this section, we aim to give an analytic proof of Corollary 2.5. Note that the generating function of $\overline{M}_{k}(n)$ is
$\sum_{n=1}^{\infty}\overline{M}_{k}(n)q^{n}=2\frac{(-q;q)_{k}}{(q;q)_{k}}\sum_{j=0}^{\infty}\frac{q^{(k+1)(k+j+1)}(-q^{k+j+2};q)_{\infty}}{(q^{k+j+1};q)_{\infty}},$
so it remains to show
$\begin{split}&\quad\frac{(-q;q)_{k-1}}{(q;q)_{k-1}}\sum_{j=0}^{\infty}\frac{q^{k(k+j)}(-q^{k+j+1};q)_{\infty}}{(q^{k+j};q)_{\infty}}-\frac{(-q;q)_{k}}{(q;q)_{k}}\sum_{j=0}^{\infty}\frac{q^{(k+1)(k+j+1)}(-q^{k+j+2};q)_{\infty}}{(q^{k+j+1};q)_{\infty}}\\ &=\frac{(-q;q)_{k}}{(q;q)_{k-1}}\sum_{j=0}^{\infty}\frac{q^{k(k+j)}(-q^{k+j+1};q)_{\infty}}{(q^{k+j+1};q)_{\infty}}.\end{split}$ (5.1)
Dividing both sides of (5.1) by $\frac{(-q;q)_{k-1}}{(q;q)_{k}}$, we find that it is equivalent to
$\begin{split}&\quad(1-q^{k})\sum_{j=0}^{\infty}\frac{q^{k(k+j)}(-q^{k+j+1};q)_{\infty}}{(q^{k+j};q)_{\infty}}-(1+q^{k})\sum_{j=0}^{\infty}\frac{q^{(k+1)(k+j+1)}(-q^{k+j+2};q)_{\infty}}{(q^{k+j+1};q)_{\infty}}\\ &=(1-q^{2k})\sum_{j=0}^{\infty}\frac{q^{k(k+j)}(-q^{k+j+1};q)_{\infty}}{(q^{k+j+1};q)_{\infty}}.\end{split}$ (5.2)
Then, we proceed to show (5.2).
Proof of (5.2). Note that
$\quad(1+q^{k})q^{(k+1)(k+j+1)}$
$=q^{(k+1)(k+j+1)}+q^{(k+1)(k+j+1)+k}$
$=q^{k(k+j+1)}(q^{k+j+1}-1+1)+q^{k(k+j+2)}(q^{k+j+1}+1-1)$
$=q^{k(k+j+1)}(1-q^{k})-q^{k(k+j+1)}(1-q^{k+j+1})+q^{k(k+j+2)}(1+q^{k+j+1}),$
so we get
$\quad(1+q^{k})\sum_{j=0}^{\infty}\frac{q^{(k+1)(k+j+1)}(-q^{k+j+2};q)_{\infty}}{(q^{k+j+1};q)_{\infty}}$
$=\sum_{j=0}^{\infty}\left((1-q^{k})\frac{q^{k(k+j+1)}(-q^{k+j+2};q)_{\infty}}{(q^{k+j+1};q)_{\infty}}-\frac{q^{k(k+j+1)}(-q^{k+j+2};q)_{\infty}}{(q^{k+j+2};q)_{\infty}}\right$.
$\qquad\qquad\qquad\left.+\frac{q^{k(k+j+2)}(-q^{k+j+1};q)_{\infty}}{(q^{k+j+1};q)_{\infty}}\right)$
$=(1-q^{k})\sum_{j=1}^{\infty}\frac{q^{k(k+j)}(-q^{k+j+1};q)_{\infty}}{(q^{k+j};q)_{\infty}}-\sum_{j=1}^{\infty}\frac{q^{k(k+j)}(-q^{k+j+1};q)_{\infty}}{(q^{k+j+1};q)_{\infty}}$
$\qquad\qquad\qquad+\sum_{j=0}^{\infty}\frac{q^{k(k+j+2)}(-q^{k+j+1};q)_{\infty}}{(q^{k+j+1};q)_{\infty}}.$
Then, we have
$\quad(1-q^{k})\sum_{j=0}^{\infty}\frac{q^{k(k+j)}(-q^{k+j+1};q)_{\infty}}{(q^{k+j};q)_{\infty}}-(1+q^{k})\sum_{j=0}^{\infty}\frac{q^{(k+1)(k+j+1)}(-q^{k+j+2};q)_{\infty}}{(q^{k+j+1};q)_{\infty}}$
$=(1-q^{k})\frac{q^{k^{2}}(-q^{k+1};q)_{\infty}}{(q^{k};q)_{\infty}}+\sum_{j=1}^{\infty}\frac{q^{k(k+j)}(-q^{k+j+1};q)_{\infty}}{(q^{k+j+1};q)_{\infty}}-\sum_{j=0}^{\infty}\frac{q^{k(k+j+2)}(-q^{k+j+1};q)_{\infty}}{(q^{k+j+1};q)_{\infty}}$
$=\sum_{j=0}^{\infty}\frac{q^{k(k+j)}(-q^{k+j+1};q)_{\infty}}{(q^{k+j+1};q)_{\infty}}-\sum_{j=0}^{\infty}\frac{q^{k(k+j+2)}(-q^{k+j+1};q)_{\infty}}{(q^{k+j+1};q)_{\infty}}$
$=(1-q^{2k})\sum_{j=0}^{\infty}\frac{q^{k(k+j)}(-q^{k+j+1};q)_{\infty}}{(q^{k+j+1};q)_{\infty}}.$
We arrive at (5.2). This completes the proof.